Assume that the rate of accretion of material to a spherical planet is given by

dV/dt = A+BV

Where V is volume, t is time and A and B are constants. Neglecting compression of the material that accretes with a density p:

(a) Write an expression for the radius of the planet with time if B=0.

(b) Write an expression for the radius of the planet with time if A=0 and at time t=0 the volume of the planet is V0.

(c) For case (b) write an expression for the average temperature of the planet with time if the specific heat of the material forming the planet is CP.


1
Expert's answer
2021-01-19T11:59:20-0500

a)

"\frac{dV}{dt}=A\\V=\frac{4\pi}{3}r^3=\frac{4\pi}{3}r^3_0+At\\r=\left(r^3_0+\frac{3}{4\pi}At\right)^{\frac{1}{3}}"

b)


"\ln{\frac{r^3}{r_0^3}}=B\\r=r_0\left(e^{Bt}\right)^{\frac{1}{3}}"

c)


"1+\alpha(T-T_0)=\left(e^{Bt}\right)^{\frac{1}{3}}"